Single plantwide and Multiple production department factory overhead rate methods and product cost distortion Instructions Single Plantwide Method Multiple Production Department Method Final Questions Instructions The management of Firebolt Industries Inc. manufactures gasoline and diesel engines through two production departments, Fabrication and Assembly. Management needs accurate product cost information in order to guide product strategy. Presently, the company uses a single plantwide factory overhead rate for allocating factory overhead to the two products. However, management is considering the multiple production department factory overhead rate method. The following factory overhead was budgeted for Firebolt: 1 Fabrication Department factory overhead 2 Assembly Department factory overhead 3 Total 636,650.00 43,000.00 879,650.00Single plantwide and Multiple production department factory overhead rate methods and product cost distortion Instructions Single Plantwide Method Multiple Production Department Method Final Questions Instructions Multiple Production Department Method b. Determine the per-unit factory overhead allocated to the gasoline and diesel engines under the muitiple production department factory overhead rate method, using direct labor hours as the activity base for each department. If required, round all per-unit answers to the nearest cent Gasoline engine $ Diesel engine $ per unit per unitFinal Questions c. (1) Recommend to management a product costing approach, based on your analyses in (a) and (b). (2) Give a reason for your answer Management is indifferent, since either method yields the same result O Management should continue to use the single plantwide overhead rate method O Management should change to the multiple production department factory overhead rate method. In this case, the single plantwide method causes cost distortion, so the multiple production department method should be used. O In this case, the multiple production department method causes cost distortion, so the single plantwide method should be used. In this case, the factory overhead rates for each product are the same under either method; therefore, the company should choose single plantwide method since its easier to implement.Instructions The management of Firebolt Industries Inc. manufactures gasoline and diesel engines through two production departments, Fabrication and Assembly. Management needs accurate product cost information in order to guide product strategy. Presently, the company uses a single plantwide factory overhead rate for allocating factory overhead to the two products. However, management is considering the multiple production department factory overhead rate method. The following factory overhead was budgeted for Firebolt: 1 Fabrication Department factory overhead 2Assembly Department factory overhead 3 Total 636,650.00 243,000.00 $879,650.00Direct labor hours were estimated as follows Fabrication Department 5,350 hours Assembly Department 5,400 Total 10,750 hours In addition, the direct labor hours (dlh) used to produce a unit of each product in each department were determined from engineering records, as follows: Production Departments Gasoline Engine Diesel Engine Fabrication Department 2.9 dlh Assembly Department 1.9 Direct labor hours per unit 4.8 dih 1.9 dlh 2.9 4.8 dlhRequired a. Determine the per-unit factory overhead allocated to the gasoline and diesel engines under the single plantwide factory overhead rate method using direct labor hours as the activity base. b. Determine the per-unit factory overhead allocated to the gasoline and diesel engines under the multiple production department factory overhead rate method, using direct labor hours as the activity base for each department. c. (1) Recommend to management a product costing approach, based on your analyses in (a) and (b). (2) Give a reason for your answer. If required, round all per-unit answers to the nearest cent.Shaded cells have feedback. X Multiple Production Department Method b. Determine the per unit factory overhead allocated to the gasoline and diesel engines under the multiple production department actor each deparmnt required, round all per-unit answers to the nearest cent. overhead rate m to using recta rh urs she act y base Gasoline engine Diesel engine per unit per unit N Points: 0/2 Feedback Check My Work Department rate Department Overhead Labor hours Product allocation Labor hours per product x Department rate; Add both department allocations together to obtain a total per product.
Answer #1
A
Overhead rate 81.83 =879650/10750
Gasoline engine 392.78 per unit =4.8*81.83
Diesel engine 392.78 per unit =4.8*81.83
B
Overhead rate:
Fabrication 119 =636650/5350
Assembly 45 =243000/5400
Gasoline engine 430.60 per unit =(2.9*119)+(1.9*45)
Diesel engine 356.60 per unit =(1.9*119)+(2.9*45)
C
1
Management should change to the multiple production department factory overhead rate method.
2
In this case, the single plantwide method causes cost distortion, so the multiple production department method should be used.